OpenStudy (anonymous):

Solve by the elimination method 0.3x-0.2y=4 0.4x+0.5y=1 Simplified answer. In an ordered pair. Use integers or fractions for any numbers in the expression

OpenStudy (anonymous):

multiply the first equation by 5 and the second by 2. 1.5x-y=20 .8x+y=2 Add them to get 2.3x=22 x=220/23. Plug this back in to get the y value for your ordered pair

OpenStudy (anonymous):

?